OpenStudy (anonymous):

Write the equation y - 2 = 2(x - 3) in standard form. A. y - 2x = 4 B. 2x - y = -6 C. 2x + y = 6 D. 2x - y = 4

OpenStudy (jhannybean):

\[y-2=2x-6\]\[y-2x=-6+2\]\[y-2x= -4\]

OpenStudy (jhannybean):

and then you can change it around i believe.